A 40.0 gram-sample of methanol, CH4O is mixed with 60.0 grams of ethanol, C2H60. What is the mole fraction of the methanol?


1
Expert's answer
2022-03-17T14:57:30-0400

Solution.

"m(CH_4O)=40.0g;"

"m(C_2H_6O)=60.0g;"

"M(CH_4O)=32 g\/mol;"

"M(C_2H_6O)=46 g\/mol;"

"\nu(CH_4O)=\dfrac{40.0}{32}=1.25mol;"

"\nu(C_2H_6O)=\dfrac{60.0}{46}=1.3mol;"

"X_i=\dfrac{1.25}{1.25+1.3}=0.49;"

Answer: "X_i=0.49."